# Generating Exceptional Behavior Tests with Reasoning Augmented Large Language Models

### v2

@@ III-A4 Connecting EBTs to relevant non-EBTs
We add the relevant non-EBTs to the prompt (lines 17 to 24 in Algorithm 1), to encourage the LLM to reason about the procedures to set up the object under test and the condition under which the exception will be triggered grounding the existing non-EBTs. Additionally, we believe the non-EBTs in the same repository will promote the consistency between the generated code and the existing code in terms of the both format and coding conventions. Given an MUT, we use two approaches to retrieve the relevant non-EBTs: non-EBTs that directly invoke the same MUT and non-EBTs that are already present in the destination test file (line 23 in Algorithm 1). If there is no relevant non-EBT retrieved, this part of the prompt is left empty.
@@ III-A5 Instruction fine-tuning
We use CodeLlama-Instruct-7B [39], an open-source foundation model designed for code generation and instruction following, as the foundation model of exLóng. CodeLlama is pretrained on auto-regressive and infilling objectives, enabling tasks like code completion and document generation. We instruction fine-tune CodeLlama on our collected training corpus of instruction data $\mathcal{C}=\{(\mathtt{p},\mathtt{t}_{\mathtt{eb}})\}$, with the novel instructions shown in Figure 2 ( 4). Given the instruction that includes the collected context, the model is expected to produce the EBT: $\mathtt{LLM}(\mathtt{p})=\mathtt{t}_{\mathtt{eb}}$.
We instruction fine-tune the CodeLlama model using the parameter-efficient Low-Rank Adaptation (LoRA) technique [21]. Rather than updating the entire set of parameters in the LLM, LoRA injects trainable low-rank matrices into each layer of the model. This approach dramatically reduces the number of trainable parameters and the amount of required computational resources.
@@ III-B Inference
Algorithm 4 Prepare the pool stack traces from non-EBTs and assemble the prompt for evaluating exLóng.
@@
15:inputs: $\mathtt{m_{ut}}$, $\mathtt{s}$, $\mathtt{d}$ task inputs selected by developers or inferred    $\mathtt{T}_{\mathtt{neb}}$ existing non-EBTs
16:outputs: $\mathtt{p}$ the prompt to give to the LLM
17:procedure AssemblePrompt($\mathtt{m_{ut}}$, $\mathtt{s}$, $\mathtt{d}$)
18:    $\mathtt{R}\leftarrow\emptyset$
@@
20:    for $\mathtt{q}$in $\mathcal{Q}$ do
21:         if $\mathtt{m_{ut}}$in $\mathtt{q}$.$\mathtt{r}$land $\mathtt{q}$.$\mathtt{s}$ == $\mathtt{s}$ then
22:             $\mathtt{R}\leftarrow\mathtt{R}\ \cup$$\mathtt{q}$.$\mathtt{r}$}
23:             if $\mathtt{m_{ut}}$ == $\mathtt{q}$.$\mathtt{r}$[0].method then
24:                 $\hat{\mathtt{T}}_{\mathtt{neb}}$leftarrow $\hat{\mathtt{T}}_{\mathtt{neb}}$cup $\mathtt{q}$.$\mathtt{t}_{\mathtt{neb}}$}
25:    if $\mathtt{R}\neq\emptyset$ then
26:         $\mathtt{r}$leftarrow RandomSelect($\mathtt{R}$)
27:         $\mathtt{g}$ = ExtractGuardExp($\mathtt{r}$)
28:         $\hat{\mathtt{T}}_{\mathtt{neb}}$leftarrow $\hat{\mathtt{T}}_{\mathtt{neb}}$cup $\{\mathtt{t}_{\mathtt{neb}}|\text{GetFile}(\mathtt{t}_{\mathtt{neb}})\text{==}d\}$
29:         return ($\mathtt{m_{ut}}$, $\mathtt{s}$, $\mathtt{d}$, $\mathtt{r}$, $\mathtt{g}$, $\hat{\mathtt{T}}_{\mathtt{neb}}$)
The inference workflow of exLóng is different from its training workflow in that we cannot rely on executing EBTs to collect the context (e.g., stack trace), as our goal is to generate those EBTs. Instead, exLóng reasons about the context based on task inputs (MUT, target throw statement, and destination test file) and leveraging existing non-EBTs for the system under test. Algorithm 4 describes the key steps in preparing the inference prompt of exLóng.
@@ III-B1 Collecting non-EBTs’ stack traces to reach potential target throw statements
We first prepare a set of stack traces, from the execution of non-EBTs, that can reach potential target throw statements in the repository (lines 4 to 12 in Algorithm 4). This only needs to be done once per repository. Specifically, we first instrument all methods that have throw statements to log the current stack trace (using $\mathtt{Thread.currentThread().getStackTrace();}$) upon invoking those methods (line 5 in Algorithm 4). Then, we execute all non-EBTs and collect the logged stack traces (lines 7 to 11 in Algorithm 4). The execution of each non-EBT may generate multiple stack traces, as it may cover multiple methods with throw statements.
@@ III-B2 Selecting task inputs
Next, we select the task inputs (MUT, target throw statement, and destination test file), which can be specified by the developer or inferred by heuristics depending on the use case that exLóng is targeting:
@@
  • Machine-oriented use case: given a repository, exLóng locates all throw statements and generates one EBT for each of them. For a target throw statement, the MUT is the method containing the throw statement, and the destination test file is selected based on (a) file name matching, and (b) test coverage analysis. Specifically, similar to prior work [36], given a code file named FNM, we search for test file named FNMTest or TestFNM. If there is no result based on file name matching, we run the existing non-EBTs to find any existing test class that cover the MUT or the class of MUT. If there is again no result based on test coverage analysis, exLóng will not generate an EBT for this target throw statement (and move to the next one).
The selected task inputs are then used to assemble the prompt for exLóng (line 15 in Algorithm 4).
@@ III-B3 Assembling the prompt
We first need to find stack traces from the set of non-EBTs’ stack traces that match the given MUT and target throw statement (line 21 in Algorithm 4). If multiple matching stack traces are found, we randomly select one (line 26 in Algorithm 4). Given the stack trace, we use the same algorithm in Section III-A3 to compute the corresponding guard expression (line 27 in Algorithm 4). The relevant non-EBTs are selected using the similar criteria as Section III-A4: have the same MUT (line 24 in Algorithm 4) or be in the same destination test file (line 28 in Algorithm 4). However, if no matching stack trace is found, exLóng will not generate an EBT for the given inputs.
@@ IV-B Dataset Statistics
The dataset is randomly split by projects into training (Train), validation (Valid), and evaluation (Eval) sets, where the training set is used to instruction fine-tune exLóng, the validation set is used for early-stopping the training process and guiding our design decision of exLóng, and the evaluation set is used for evaluating the performance of exLóng and baselines.
Table II presents the statistics of the evaluation data for developer-oriented use case. Note that this is a subset of the last row from Table I. Under developer-oriented use case, we benchmark exLóng on the subset of 434 data points for which we are able to extract stack traces.
Table III presents the statistics of the evaluation data for machine-oriented use case. Note that this is a subset of the last row from Table I. For machine-oriented use case, we evaluate on 649 data points as we filter the data for which we were not able to locate the destination test file with our designed heuristics (Section III-B2).
@@ V Evaluation Design
We assess the performance of exLóng by answering the following research questions:
RQ1: How does exLóng perform under the developer-oriented use case compared with the state-of-the-art models?
RQ2: How much do stack traces and guard expressions help exLóng in generating EBTs?
RQ3: How much does the selection of non-EBTs help exLóng in generating EBTs?
RQ4: How does the size of the underlying LLM model impact our findings?
RQ5: How does exLóng perform under the machine-oriented use case compared with analysis-based test generation tools?
We next describe metrics used to compare models and tools (Section V-A) and then describe the baselines used in our comparison (Section V-B) and hardware used for experiments (Section  V-C). We answer all research questions in Section VI.
@@ V-A1 Developer-Oriented Use Case
The similarity metrics only capture the surface-level similarity between the prediction against an existing EBT. However, such surface metrics do not adequately capture the functional validity of the generated EBT (e.g., whether the code can be compiled or executed), especially since the developer-written EBTs may not be the only correct implementation to cover a specific target throw statement. Thus, we additionally include the following functional-correctness metrics:
Compilable%: percentage of the generated EBTs that can be compiled.
@@
Runnable%: percentage of EBTs that check the specified exception type, and can be compiled and executed without any error.
ThrowCov%: out of all developer-specified target throw statements (Table II), the percentage of target throw statements with successfully generated EBTs, i.e., compilable, runnable, and checking the specified exception type.
@@ V-B1 Learning-based tools
We compare exLóng with one of the strongest foundation models and one LLM that is specifically pretrained to generate tests.
GPT3.5: We instruct GPT3.5 [5, 31] to write EBTs by first providing one random example from the training data. Namely, one prompt and the corresponding ground-truth EBT. The prompt we use to query GPT3.5 includes the MUT, the target exception type to test, the method containing the target throw statement, one relevant non-EBT, and the destination test file. We sample a single EBT from the output.
@@ V-B2 Automatic test generation tools
In machine-oriented use case, we compare exLóng with two widely-used analysis-based test generation tools.
@@ V-C Hardware
We fine-tune exLóng on 4 Nvidia A100 GPUs, 2 AMD Milan 7413 @ 2.65 GHz. We run Randoop and EvoSuite on a machine with Intel Core i7-11700K @ 3.60GHz (8 cores, 16 threads) CPU, 64 GB RAM, Ubuntu 20.04, Java 8, and Maven 3.8.6. All random seeds are set to 42.
@@ VI Results
CAT-LM 53.96 60.41 80.00 10.14 71.20 100.00 40.32 32.26
exLóng 64.17 68.60 85.78 19.82 82.49 100.00 67.97 58.53
exLóng-sample 69.60 73.40 88.77 18.89 87.10 100.00 75.12 66.82
@@
CAT-LM 38.00 49.11 69.55 1.38 70.05 100.00 25.35 17.05
exLóng 47.84 55.89 80.29 2.07 82.95 100.00 69.59 58.99
exLóng-sample 51.58 59.76 83.23 1.84 89.63 100.00 79.49 68.66
@@ VI-A RQ1: Developer-Oriented Use Case
To answer RQ1, we compare the EBTs generated by exLóng with developer-written tests. The results of exLóng and baselines are shown in tables IV and V. Table IV presents the results when we inform LLMs the method name of the target EBT while in Table V we do not. (We describe the last row in these tables in a later subsection.) Note that we add the same random-selected relevant non-EBT to the prompt for all the models.
exLóng outperforms all the baselines on both similarity-based metrics (left side in tables) and functional-correctness metrics (right side in tables). exLóng achieves higher performance than baselines for both generating executable EBTs (Runnable%) and EBTs that cover the target throw statements (ThrowCov%). In Table IV, we can see that exLóng outperforms GPT3.5 by 10.9% and 20.9% on Runnableand ThrowCov%, respectively. Similarly, we can see that exLóng outperforms CAT-LM by 68.6% and 57.53% on Runnableand ThrowCov%, respectively. This highlights the benefit of the stack traces and guard expressions extracted via program analysis, and exLóng’s capability of reasoning about them.
Comparing functional-correctness metrics in Table IV and Table V, performance of both GPT3.5 and CAT-LM declines significantly when the EBT method name is omitted. This result aligns with expectations, as the method name frequently implies the conditions under which the exception is supposed to be thrown, e.g., “$\mathtt{should\_fail\_if\_time\_provider\_is\_null}$”. In contrast, exLóng demonstrates robustness regarding the inclusion or exclusion of the test method name, maintaining consistent performance on functional-correctness metrics. This further emphasizes the reasoning ability of exLóng on the stack traces and guard expressions.
@@ VI-B RQ2: Ablation Study of Stack Traces and Conditions
To evaluate the contribution of the components in exLóng, we perform an ablation study. In Table VI, we show the results while including EBT’s name in the prompt. We find that ablating each component deteriorates performance especially across functional correctness metrics. Removing stack traces slightly hurts the performance of exLóng which is expected, because a guard expression is, in a way, the summary of a stack trace.
We observe a small drop in Compilebut a rather larger drop in ThrowCovif removing both stack trace and guard expression from the context. This underlines the importance of the condition information present by these two components on which exLóng reasons about when generating EBTs that cover the target throw statements. Relevant non-EBTs mostly contribute to the Compilebecause the relevant non-EBTs covers the same MUT which gives model a starting point to construct the EBTs.

### v3

@@ III-A4 Connecting EBTs to relevant non-EBTs
We add the relevant non-EBTs to the prompt (lines 17 to 24 in Algorithm 1), to encourage the LLM to reason about the procedures to set up the object under test and the condition under which the exception will be triggered grounding the existing non-EBTs. Additionally, we believe the non-EBTs in the same repository will promote the consistency between the generated code and the existing code in terms of the both format and coding conventions. Given an MUT, we use two approaches to retrieve the relevant non-EBTs: non-EBTs that directly invoke the same MUT (used first when the context window of the LLM cannot fit all relevant non-EBTs) and non-EBTs that are already present in the destination test file (line 23 in Algorithm 1). If there is no relevant non-EBT retrieved, this part of the prompt is left empty.
@@ III-A5 Instruction fine-tuning
We use CodeLlama-Instruct-7B [43], an open-source foundation model designed for code generation and instruction following, as the foundation model of exLong. CodeLlama is pretrained on auto-regressive and infilling objectives, enabling tasks like code completion and document generation. We fine-tune CodeLlama on our collected training corpus of SFT data $\mathcal{C}=\{(\mathtt{p},\mathtt{t}_{\mathtt{eb}})\}$, with the novel instructions shown in Figure 2 ( 4). Given the instruction that includes the collected context, the model is expected to produce the EBT: $\mathtt{LLM}(\mathtt{p})=\mathtt{t}_{\mathtt{eb}}$.
We fine-tune the CodeLlama model using the parameter-efficient Low-Rank Adaptation (LoRA) technique [20]. Rather than updating the entire set of parameters in the LLM, LoRA injects trainable low-rank matrices into each layer of the model. This approach dramatically reduces the number of trainable parameters and the amount of required computational resources.
@@ III-B Inference
Algorithm 4 Prepare the pool stack traces from non-EBTs and assemble the prompt for evaluating exLong.
@@
15:inputs: $\mathtt{m_{ut}}$, $\mathtt{s}$, $\mathtt{d}$ task inputs selected by developers or inferred
16:    $\mathtt{T}_{\mathtt{neb}}$ existing non-EBTs
17:outputs: $\mathtt{p}$ the prompt to give to the LLM
18:procedure AssemblePrompt($\mathtt{m_{ut}}$, $\mathtt{s}$, $\mathtt{d}$)
19:    $\mathtt{R}\leftarrow\emptyset$
@@
21:    for $\mathtt{q}$in $\mathcal{Q}$ do
22:         if $\mathtt{m_{ut}}$in $\mathtt{q}$.$\mathtt{r}$land $\mathtt{q}$.$\mathtt{s}$ == $\mathtt{s}$ then
23:             $\mathtt{R}\leftarrow\mathtt{R}\ \cup$$\mathtt{q}$.$\mathtt{r}$}
24:             if $\mathtt{m_{ut}}$ == $\mathtt{q}$.$\mathtt{r}$[0].method then
25:                 $\hat{\mathtt{T}}_{\mathtt{neb}}$leftarrow $\hat{\mathtt{T}}_{\mathtt{neb}}$cup $\mathtt{q}$.$\mathtt{t}_{\mathtt{neb}}$}
26:    if $\mathtt{R}\neq\emptyset$ then
27:         $\mathtt{r}$leftarrow RandomSelect($\mathtt{R}$)
28:         $\mathtt{g}$ = ExtractGuardExp($\mathtt{r}$)
29:         $\hat{\mathtt{T}}_{\mathtt{neb}}$leftarrow $\hat{\mathtt{T}}_{\mathtt{neb}}$cup $\{\mathtt{t}_{\mathtt{neb}}|\text{GetFile}(\mathtt{t}_{\mathtt{neb}})\text{==}d\}$
30:         return ($\mathtt{m_{ut}}$, $\mathtt{s}$, $\mathtt{d}$, $\mathtt{r}$, $\mathtt{g}$, $\hat{\mathtt{T}}_{\mathtt{neb}}$)
The inference workflow of exLong is different from its training workflow in that we cannot rely on executing EBTs to collect the context (e.g., stack trace), as our goal is to generate those EBTs. Instead, exLong reasons about the context based on task inputs (MUT, target throw statement, and destination test file) and leveraging existing non-EBTs for the system under test. Algorithm 4 describes the key steps in preparing the inference prompt of exLong.
@@ III-B1 Collecting non-EBTs’ stack traces to reach potential target throw statements
We first prepare a set of stack traces, from the execution of non-EBTs, that can reach potential target throw statements in the repository (lines 4 to 12). This only needs to be done once per repository. Specifically, we first instrument all methods that have throw statements to log the current stack trace (using $\mathtt{Thread.currentThread().getStackTrace();}$) upon invoking those methods (line 5). Then, we execute all non-EBTs and collect the logged stack traces (lines 7 to 11). The execution of each non-EBT may generate multiple stack traces, as it may cover multiple methods with throw statements.
@@ III-B2 Selecting task inputs
Next, we select the task inputs (MUT, target throw statement, and destination test file), which can be specified by the developer or inferred by heuristics depending on the use case that exLong is targeting:
@@
  • Machine-oriented use case: given a repository, exLong locates all throw statements and generates one EBT for each of them. For a target throw statement, the MUT is the method containing the throw statement, and the destination test file is selected based on (a) file name matching, and (b) test coverage analysis. Specifically, similar to prior work [40], given a code file named FNM, we search for test file named FNMTest or TestFNM. If there is no result based on file name matching, we run the existing non-EBTs to find any existing test class that cover the MUT or the class of MUT. If there is again no result based on test coverage analysis, exLong will not generate an EBT for this target throw statement (and it will move to the next one).
The selected task inputs are then used to assemble the prompt for exLong (line 15).
@@ III-B3 Assembling the prompt
We first need to find stack traces from the set of non-EBTs’ stack traces that match the given MUT and target throw statement (line 22). If multiple matching stack traces are found, we randomly select one (line 27). Given the stack trace, we use the same algorithm in Section III-A3 to compute the corresponding guard expression (line 28). The relevant non-EBTs are selected using the similar criteria as Section III-A4, i.e., having the same MUT (line 25) or in the same destination test file (line 29). However, if no matching stack trace is found, exLong will not generate an EBT for the given inputs.
@@ IV-B Dataset Statistics
The dataset is randomly split by projects into training (Train), validation (Valid), and evaluation (Eval) sets, where the training set is the SFT data used to instruction fine-tune exLong, the validation set is used for early-stopping the training process and guiding our design decision of exLong, and the evaluation set is used for evaluating the performance of exLong and baselines.
Table II presents the statistics of the evaluation data for developer-oriented use case. Note that this is a subset of the last row from Table I. Under developer-oriented use case, we benchmark exLong on the subset of 434 examples for which we are able to extract stack traces. In this paper, we focus on cases where accurate stack traces can be extracted by executing existing non-EBTs. When such non-EBTs are not available, namely the stack traces cannot be obtained, developers can first write or generate non-EBTs for the MUT with the help of other test generation tools, and then use exLong to generate EBTs.
Table III presents the statistics of the evaluation data for machine-oriented use case. Note that this is a subset of the last row from Table I. For machine-oriented use case, we evaluate on 649 examples as we filter the data for which we were not able to locate the destination test file with our designed heuristics (Section III-B2).
@@ V Evaluation Design
We assess the performance of exLong by answering the following research questions:
RQ1: How does exLong perform under the developer-oriented use case compared with the state-of-the-art models?
RQ2: How much do stack traces and guard expressions help exLong in generating EBTs?
RQ3: How much does the selection of non-EBTs help exLong in generating EBTs?
RQ4: How does exLong perform with different underlying LLM model?
RQ5: How does exLong perform under the machine-oriented use case compared with analysis-based test generation tools?
We next describe metrics used to compare models and tools (Section V-A) and then describe the baselines used in our comparison (Section V-B). We answer all research questions in Section VI.
@@ V-A1 Developer-oriented use case
The similarity metrics only capture the surface-level similarity between the prediction against an existing EBT; among them, xMatch is the most strict one as it requires perfect matches, while the others account for partial matches. However, such surface metrics do not adequately capture the functional validity of the generated EBT (e.g., whether the code can be compiled or executed), especially since the developer-written EBTs may not be the only correct implementation to cover a specific target throw statement. Thus, we additionally include the following functional-correctness metrics:
Compilable%: percentage of the generated EBTs that can be compiled. Being compilable is a basic functional requirement for the generated tests.
@@
Runnable%: percentage of EBTs that check the specified exception type, and can be compiled and executed without any error. This metric, unlike others, requires the generated EBTs to be semantically valid.
ThrowCov%: out of all developer-specified target throw statements (Table II), the percentage of target throw statements with successfully generated EBTs, i.e., compilable, runnable, and checking the specified exception type. This is the strictest metric, ensuring that the generated EBTs are semantically valid and are targeting the throw statement specified by developers.
@@ V-B1 Learning-based tools
We compare exLong with one of the strongest foundation models and one LLM that is specifically pretrained to generate tests.
GPT3.5: We instruct GPT3.5 [34] to write EBTs by first providing one random example from the training data. Namely, one prompt and the corresponding ground-truth EBT. The prompt we use to query GPT3.5 includes the MUT, the target exception type to test, the method containing the target throw statement, one relevant non-EBT, and the destination test file. We sample a single EBT from the output.
@@ V-B2 Automatic test generation tools
In machine-oriented use case, we compare exLong with two widely-used analysis-based test generation tools.
@@ V-C Hardware
We run exLong’s program analyses part, Randoop, and EvoSuite on a machine with Intel Core i7-11700K @ 3.60GHz (8 cores, 16 threads) CPU, 64 GB RAM, Ubuntu 20.04, Java 8, and Maven 3.8.6. We perform exLong’s LLM fine-tuning and generation, as well as CAT-LM on a server with 4 Nvidia A100 GPUs, 2 AMD Milan 7413 @ 2.65 GHz. We run fine-tuning and generation, for exLong and baselines, three times with different random seeds and report the average numbers across three runs.
@@ VI Results
CAT-LM 53.49 59.79 78.91 9.83 71.83 100.00 36.64 30.03
CodeLlama zero-shot 38.23 48.63 66.96 5.53 57.30 96.02 40.17 33.87
exLong 63.13 67.49 85.32 19.05 82.10 100.00 67.36 59.45
exLong-sample 70.01 74.08 90.09 15.28 93.54 100.00 81.29 71.28
@@
CAT-LM 37.25 48.47 68.18 1.23 70.83 100.00 23.96 16.44
CodeLlama zero-shot 26.93 37.38 61.88 0.15 57.22 96.91 41.32 33.49
exLong 46.66 55.36 79.76 2.07 82.26 100.00 69.89 59.83
exLong-sample 50.73 59.38 82.83 1.92 89.40 100.00 78.80 67.67
@@ VI-A RQ1: Developer-Oriented Use Case
To answer RQ1, we compare the EBTs generated by exLong with developer-written tests. The results of exLong and baselines are shown in tables IV and V. Table IV presents the results when we inform LLMs the method name of the target EBT while in Table V we do not. (We describe the last row in these tables in a later subsection.)
exLong outperforms all the baselines on both similarity-based metrics (left side in tables) and functional-correctness metrics (right side in tables). exLong achieves higher performance than baselines for both generating executable EBTs (Runnable%) and EBTs that cover the target throw statements (ThrowCov%). This highlights that exLong can generate more EBTs that can be directly adopted by developers. In Table IV, we can see that exLong outperforms GPT3.5 by 9.9% and 22.8% on Runnableand ThrowCov%, respectively. Similarly, we can see that exLong outperforms CAT-LM by 83.8% and 97.5% on Runnableand ThrowCov%, respectively. This further underlines the benefit of the stack traces and guard expressions extracted via program analysis, and exLong’s capability of reasoning about them.
[…]
public static FileWriter createFileWriter(String className, LogFilePath logFilePath, CompressionCodec codec, SecorConfig config)
throws Exception {
return createFileReaderWriterFactory(className, config).BuildFileWriter(logFilePath, codec);
((a)) The MUT to be tested.
!FileReaderWriterFactory.class.isAssignableFrom(Class.forName(className))
((b)) The guard expression for the target throw statement.
@Test(expected = IllegalArgumentException.class)
public void testFileWriterConstructorMissing() throws Exception {
ReflectionUtil.createFileWriter("MissingClass", mLogFilePath, null, mSecorConfig);
((c)) Compilable but failing EBT generated by GPT3.5
@Test(expected = IllegalArgumentException.class)
public void testFileWriterConstructorMissing() throws Exception {
ReflectionUtil.createFileWriter("java.lang.String", mLogFilePath, null, mSecorConfig);
((d)) Compilable and runnable EBT generated by exLong
[…]
Comparing functional-correctness metrics in Table IV and Table V, performance of both GPT3.5 and CAT-LM declines significantly when the EBT method name is omitted. This result aligns with expectations, as the method name frequently implies the conditions under which the exception is supposed to be thrown, e.g., “$\mathtt{should\_fail\_if\_time\_provider\_is\_null}$”. In contrast, exLong demonstrates robustness regarding the inclusion or exclusion of the test method name, maintaining consistent performance on functional-correctness metrics. This further emphasizes the reasoning ability of exLong on the stack traces and guard expressions.
@@ VI-B RQ2: Ablation Study of Stack Traces and Conditions
To evaluate the contribution of the components in exLong, we perform an ablation study. In Table VI, we show the results while including EBT’s name in the prompt. We find that ablating each component deteriorates performance especially across functional correctness metrics. Removing stack traces slightly hurts the performance of exLong in terms of functional correctness which is expected, because a guard expression is, in a way, the summary of a stack trace.
We observe a small drop in Compilablebut a rather larger drop in ThrowCovif removing both stack trace and guard expression from the context. This underlines the importance of the condition information present by these two components on which exLong reasons about when generating EBTs that cover the target throw statements. Relevant non-EBTs mostly contribute to the Compilablebecause the relevant non-EBT covers the same MUT which gives model a starting point to construct the EBTs. The ablation study underscores the importance of each component in exLong.